# Efficient Perplexity Bound and Ratio Matching in Discrete Diffusion Language Models

### v1

@@ B.3.1 Efficient Estimation of the Score in Practice
Figure 5: The conditional ratios at position $i$ over the vocabulary in the two cases. We define $b=\frac{1}{n=V}(1-e^{-\sigma_{t}})$ and $c=1-(n-1)b$.
@@
$=\sum_{j\neq{\bm{x}}^{i}_{t}}{\bm{Q}}_{t}^{tok}({\bm{x}}^{i}_{t},j)s_{j}+{\bm{Q}}_{t}^{tok}({\bm{x}}^{i}_{t},k)\left(\frac{c}{b}-1\right)s_{k}=\sum_{j\neq{\bm{x}}^{i}_{t}}\frac{1}{n}s_{j}+\frac{1}{n}\left(\frac{c}{b}-1\right)s_{k}.$ (141)
@@ B.3.2 SEDD Scaling
$=\frac{1}{n-1}\frac{e^{-\sigma_{t}}}{1-e^{-\sigma_{t}}}=\frac{1}{n-1}\frac{1}{e^{\sigma_{t}}-1}.$ (144)
@@ B.3.3 Score Estimation Through CEDD in Practice
In the case of the absorb diffusion, from Section B.3.1, we can see that we do not need the ratio when ${\bm{y}}^{i}=n$, and furthermore, we only need the ratios when ${\bm{x}}_{t}^{i}=n$. Since also the data is not masked, then ${\bm{x}}^{i}_{0}\neq n$. Therefore, the first term (the sum) in Equation (154) is 00, since $\frac{p_{t|0}({\bm{y}}^{i}\neq n|{\bm{x}}^{i}_{0}\not\in\{n,{\bm{y}}^{i}\})}{p_{t|0}(n|{\bm{x}}^{i}_{0}\not\in\{n,{\bm{y}}^{i}\})}=0$. Hence
@@
In the case of the uniform diffusion, we write $b=\frac{1}{n=V}(1-e^{-\sigma_{t}})$ and $c=1-(n-1)b$ as in Section B.3.1, and Equation (154) becomes
@@
In the case of the uniform diffusion dynamics, when $t\rightarrow 0$ then $\frac{b}{c}\rightarrow 0$ and $\frac{c}{b}\rightarrow\infty$. This does not cause any issues when sampling, however, it does negatively impact the perplexity bound. We explain informally the issue below. We fix position $i$ and assume we are at ${\bm{x}}_{t}^{i}$ and that originally we were at ${\bm{x}}_{0}^{i}$. When the time $t$ is close to 00, the model $f^{i}_{\theta}({\bm{x}}_{t},t)[{\bm{y}}^{i}]$ is typically very confident and can predict that some token ${\bm{y}}^{i}$ is ${\bm{x}}_{0}^{i}$, that is, $f^{i}_{\theta}({\bm{x}}_{t},t)[{\bm{y}}^{i}]\approx 1-\epsilon$ and $f^{i}$ is $\approx\frac{\epsilon}{V-1}$ everywhere else. Furthermore, most of the time ${\bm{x}}_{t}^{i}={\bm{x}}_{0}^{i}$. This means the score $s^{i}_{\theta}({\bm{x}}_{t},t)[{\bm{y}}^{i}]\approx\frac{c}{b}$ and is $\approx\frac{b}{c}$ everywhere else. If the prediction is correct then the loss becomes $\approx\frac{1}{V-1}(V-2)[-\frac{b}{c}\log(\frac{b}{c})+\frac{b}{c}+\frac{b}{c}(\log(\frac{b}{c})-1)]+[-\frac{c}{b}\log(\frac{c}{b})+\frac{c}{b}+\frac{c}{b}(\log(\frac{c}{b})-1)]\approx 0$. If it misses however, then we have $\approx\frac{1}{V-1}(V-3)[-\frac{b}{c}\log(\frac{b}{c})+\frac{b}{c}+\frac{b}{c}(\log(\frac{b}{c})-1)]+[-\frac{b}{c}\log(\frac{c}{b})+\frac{c}{b}+\frac{b}{c}(\log(\frac{b}{c})-1)]+[-\frac{c}{b}\log(\frac{b}{c})+\frac{b}{c}+\frac{c}{b}(\log(\frac{c}{b})-1)]\approx 2\frac{c}{b}\log\frac{c}{b}$. Considering the large magnitude of $\frac{c}{b}$ when $t\approx 0$, this loss is extremely punitive. SEDD does not suffer from this issue, as the conditional ratios $\frac{c}{b}$ are implicitly learned by the network, which due to its limited flexibility likely acts as a regularizer, not allowing the values of the ratios to rise steeply as $t\rightarrow 0$. Motivated by this, in the uniform case, we rescale $\sigma_{t}<0.0015$, by setting $\sigma_{t}=0.0015$. This significantly reduces the magnitude of $\frac{c}{b}$ as $t\rightarrow 0$. We highlight that only the $\sigma_{t}$ that is used to calculate $\frac{c}{b}$ and $\frac{b}{c}$ are scaled while the $\sigma_{t}$ that is fed to the model is not touched. Naturally, for the sake of rigor, this is also the model we employ to generate samples. Finally, we note that this problem also appears in the case of the roulette diffusion dynamics, as when time is close to 00 most tokens are unmasked. The same strategy is applied, as before, only to ratios $\frac{c}{b}$ and $\frac{b}{c}$, by rescaling $\sigma_{t}$ when $\sigma_{t}<0.5$ as follows: $\sigma^{scaled}_{t}=\log(1.1\sigma_{t}+1.1)$.
@@ B.3.4 Analytic Sampling in Practice
Since we have an analytic expression of the matrix exponential $e^{\sigma_{t}^{\Delta t}{\bm{Q}}^{tok}}$, we can easily derive the expression of $e^{-\sigma_{t}^{\Delta t}{\bm{Q}}^{tok}}$, by simply substituting $\sigma_{t}^{\Delta t}$ with $-\sigma_{t}^{\Delta t}$ in each entry. One can efficiently calculate the sum above by following the strategy of (Lou et al., 2024) as below.
In the case of the absorb diffusion, $e^{-\sigma_{t}^{\Delta t}{\bm{Q}}^{tok}}(j\neq n,j\neq n)=1-\bar{a}$, $e^{-\sigma_{t}^{\Delta t}{\bm{Q}}^{tok}}(n,n)=1$, $e^{-\sigma_{t}^{\Delta t}{\bm{Q}}^{tok}}(n,j\neq n)=\bar{a}$, with all other entries being 00, where $\bar{a}=1-e^{\sigma_{t}^{\Delta t}}$. As such if ${\bm{x}}_{t-\epsilon}\neq n$
@@
$=\bar{a}\sum_{{\bm{y}}^{i}=1}^{n}s^{i}_{\theta}({\bm{x}}_{t},t)[{\bm{y}}^{i}]+(1-\bar{a})s^{i}_{\theta}({\bm{x}}_{t},t)[n]=\bar{a}S^{i}_{\theta}({\bm{x}}_{t},t)+(1-\bar{a})s^{i}_{\theta}({\bm{x}}_{t},t)[n],$ (167)
@@
$=\frac{s^{i}_{\theta}({\bm{x}}_{t},t)[{\bm{x}}_{t-\epsilon}^{i}]}{e^{-\sigma_{t}^{\Delta t}}}+\frac{e^{-\sigma_{t}^{\Delta t}}-1}{ne^{-\sigma_{t}^{\Delta t}}}S^{i}_{\theta}({\bm{x}}_{t},t).$ (169)
@@ C.1 Generated text
boat men: drift, and boom. Built, three 9th sailors found themselves losing an average of 12 inches of the underwater 50 feet in the delta base of S. M. Gielker Jr, U.S. captain (and pilot for the 11th Brigade), sailed deep into the canal on a watchdeck with his bow of thumb and lost at his anchor near shore.
@@ D.1 Discrete-Time Markov Chains Over Finite-State Spaces
where ${\bm{P}}(x,y)=p_{t|t-1}(X_{t}=x|X_{t-1}=y)$. Given an initial probability distribution ${\bm{s}}_{0}$ over states, the equation above fully determines the evolution of the probability over states with respect to time. If it is known that the state at time $t-1$ is $y$, then ${\bm{s}}_{t}$ is simply column $y$ of ${\bm{P}}$. This implies that the sum of the elements of each column $y$ of ${\bm{P}}$ is one.
@@ D.2 Continuous-Time Markov Chains Over Finite-State Spaces
Given an initial probability distribution ${\bm{s}}_{0}$ over states, the equation above fully determines the evolution of the probability over states with respect to time. Indeed, the distribution over states at time $t$ is the solution of the linear ODE in (181): ${\bm{s}}_{t}=e^{t{\bm{Q}}}{\bm{s}}_{0}$. Therefore, if it is known that the state at time 00 is $j$, then ${\bm{s}}_{t}$ is simply column $j$ of $e^{t{\bm{Q}}}$. Naturally, one can rescale the time variable such that $t=\sigma(t)$, where $\sigma$ is monotonically increasing, $\sigma(0)=0$ and $\lim_{t\rightarrow 1}\sigma(t)=T$, so that for ${\bm{z}}_{0}:={\bm{s}}_{0}$, we get $e^{t{\bm{Q}}}{\bm{s}}_{0}=e^{\sigma(t){\bm{Q}}}{\bm{z}}_{0}={\bm{z}}_{t}$, and
@@
Matrices ${\bm{Q}}_{t}=\sigma^{{}^{\prime}}(t)({\bm{P}}-{\bm{I}})$ clearly satisfy the properties of transition-rate matrices (Suhov & Kelbert, 2008). Matrices ${\bm{Q}}_{t}$ are chosen such that: a) the matrix exponential $e^{\sigma(t){\bm{Q}}}$ is easy to calculate, which is essential as $p_{t|0}(x|y)=e^{\sigma(t){\bm{Q}}}(x,y)$; and b) ${\bm{z}}_{1}$ is an easy reference distribution to sample from (Austin et al., 2021; Campbell et al., 2022).

### v2

@@ B.3.1 Efficient Estimation of the Score in Practice
Figure 5: The conditional ratios at position $i$ over the vocabulary in the two cases. We define $b=\frac{1}{n=V}(1-e^{-\sigma_{t}})$ and c=1(n1)bc=1-(n-1)b.
@@
$\sum_{j\neq{\bm{x}}^{i}_{t}}{\bm{Q}}_{t}^{tok}({\bm{x}}^{i}_{t},j)s_{j}+{\bm{Q}}_{t}^{tok}({\bm{x}}^{i}_{t},k)\left(\frac{c}{b}-1\right)s_{k}=\sum_{j\neq{\bm{x}}^{i}_{t}}\frac{1}{n}s_{j}+\frac{1}{n}\left(\frac{c}{b}-1\right)s_{k}.$ (141)
@@ B.3.2 SEDD Scaling
$\frac{1}{n-1}\frac{e^{-\sigma_{t}}}{1-e^{-\sigma_{t}}}=\frac{1}{n-1}\frac{1}{e^{\sigma_{t}}-1}.$ (144)
@@ B.3.3 Score Estimation Through CEDD in Practice
In the case of the absorb diffusion, from Section B.3.1, we can see that we do not need the ratio when ${\bm{y}}^{i}=n$, and furthermore, we only need the ratios when ${\bm{x}}_{t}^{i}=n$. Since also the data is not masked, then ${\bm{x}}^{i}_{0}\neq n$. Therefore, the first term (the sum) in Equation (154) is 0, since $\frac{p_{t|0}({\bm{y}}^{i}\neq n|{\bm{x}}^{i}_{0}\not\in\{n,{\bm{y}}^{i}\})}{p_{t|0}(n|{\bm{x}}^{i}_{0}\not\in\{n,{\bm{y}}^{i}\})}=0$. Hence
@@
In the case of the uniform diffusion, we write $b=\frac{1}{n=V}(1-e^{-\sigma_{t}})$ and c=1(n1)bc=1-(n-1)b as in Section B.3.1, and Equation (154) becomes
@@
In the case of the uniform diffusion dynamics, when $t\rightarrow 0$ then $\frac{b}{c}\rightarrow 0$ and $\frac{c}{b}\rightarrow\infty$. This does not cause any issues when sampling, however, it does negatively impact the perplexity bound. We explain informally the issue below. We fix position $i$ and assume we are at ${\bm{x}}_{t}^{i}$ and that originally we were at ${\bm{x}}_{0}^{i}$. When the time $t$ is close to 0, the model $f^{i}_{\theta}({\bm{x}}_{t},t)[{\bm{y}}^{i}]$ is typically very confident and can predict that some token ${\bm{y}}^{i}$ is ${\bm{x}}_{0}^{i}$, that is, $f^{i}_{\theta}({\bm{x}}_{t},t)[{\bm{y}}^{i}]\approx 1-\epsilon$ and $f^{i}$ is $\approx\frac{\epsilon}{V-1}$ everywhere else. Furthermore, most of the time ${\bm{x}}_{t}^{i}={\bm{x}}_{0}^{i}$. This means the score $s^{i}_{\theta}({\bm{x}}_{t},t)[{\bm{y}}^{i}]\approx\frac{c}{b}$ and is $\approx\frac{b}{c}$ everywhere else. If the prediction is correct then the loss becomes $\approx\frac{1}{V-1}(V-2)[-\frac{b}{c}\log(\frac{b}{c})+\frac{b}{c}+\frac{b}{c}(\log(\frac{b}{c})-1)]+[-\frac{c}{b}\log(\frac{c}{b})+\frac{c}{b}+\frac{c}{b}(\log(\frac{c}{b})-1)]\approx 0$. If it misses however, then we have $\approx\frac{1}{V-1}(V-3)[-\frac{b}{c}\log(\frac{b}{c})+\frac{b}{c}+\frac{b}{c}(\log(\frac{b}{c})-1)]+[-\frac{b}{c}\log(\frac{c}{b})+\frac{c}{b}+\frac{b}{c}(\log(\frac{b}{c})-1)]+[-\frac{c}{b}\log(\frac{b}{c})+\frac{b}{c}+\frac{c}{b}(\log(\frac{c}{b})-1)]\approx 2\frac{c}{b}\log\frac{c}{b}$. Considering the large magnitude of $\frac{c}{b}$ when $t\approx 0$, this loss is extremely punitive. SEDD does not suffer from this issue, as the conditional ratios $\frac{c}{b}$ are implicitly learned by the network, which due to its limited flexibility likely acts as a regularizer, not allowing the values of the ratios to rise steeply as $t\rightarrow 0$. Motivated by this, in the uniform case, we rescale $\sigma_{t}<0.0015$, by setting $\sigma_{t}=0.0015$. This significantly reduces the magnitude of $\frac{c}{b}$ as $t\rightarrow 0$. We highlight that only the $\sigma_{t}$ that is used to calculate $\frac{c}{b}$ and $\frac{b}{c}$ are scaled while the $\sigma_{t}$ that is fed to the model is not touched. Naturally, for the sake of rigor, this is also the model we employ to generate samples. Finally, we note that this problem also appears in the case of the roulette diffusion dynamics, as when time is close to 0 most tokens are unmasked. The same strategy is applied, as before, only to ratios $\frac{c}{b}$ and $\frac{b}{c}$, by rescaling $\sigma_{t}$ when $\sigma_{t}<0.5$ as follows: $\sigma^{scaled}_{t}=\log(1.1\sigma_{t}+1.1)$.
@@ B.3.4 Analytic Sampling in Practice
Since we have an analytic expression of the matrix exponential $e^{\sigma_{t}^{\Delta t}{\bm{Q}}^{tok}}$, we can easily derive the expression of $e^{-\sigma_{t}^{\Delta t}{\bm{Q}}^{tok}}$, by simply substituting $\sigma_{t}^{\Delta t}$ with $\sigma_{t}^{\Delta t}$ in each entry. One can efficiently calculate the sum above by following the strategy of (Lou et al., 2024) as below.
In the case of the absorb diffusion, $e^{-\sigma_{t}^{\Delta t}{\bm{Q}}^{tok}}(j\neq n,j\neq n)=1-\bar{a}$, $e^{-\sigma_{t}^{\Delta t}{\bm{Q}}^{tok}}(n,n)=1$, $e^{-\sigma_{t}^{\Delta t}{\bm{Q}}^{tok}}(n,j\neq n)=\bar{a}$, with all other entries being 0, where $\bar{a}=1-e^{\sigma_{t}^{\Delta t}}$. As such if ${\bm{x}}_{t-\epsilon}\neq n$
@@
$\bar{a}\sum_{{\bm{y}}^{i}=1}^{n}s^{i}_{\theta}({\bm{x}}_{t},t)[{\bm{y}}^{i}]+(1-\bar{a})s^{i}_{\theta}({\bm{x}}_{t},t)[n]=\bar{a}S^{i}_{\theta}({\bm{x}}_{t},t)+(1-\bar{a})s^{i}_{\theta}({\bm{x}}_{t},t)[n],$ (167)
@@
$\frac{s^{i}_{\theta}({\bm{x}}_{t},t)[{\bm{x}}_{t-\epsilon}^{i}]}{e^{-\sigma_{t}^{\Delta t}}}+\frac{e^{-\sigma_{t}^{\Delta t}}-1}{ne^{-\sigma_{t}^{\Delta t}}}S^{i}_{\theta}({\bm{x}}_{t},t).$ (169)
@@ C.1 Generated text
boat men: drift, and boom. Built, three 9th sailors found themselves losing an average of 12 inches of the underwater 50 feet in the delta base of S. M. Gielker Jr, U.S. captain (and pilot for the $1$th Brigade), sailed deep into the canal on a watchdeck with his bow of thumb and lost at his anchor near shore.
@@ D.1 Discrete-Time Markov Chains Over Finite-State Spaces
where ${\bm{P}}(x,y)=p_{t|t-1}(X_{t}=x|X_{t-1}=y)$. Given an initial probability distribution ${\bm{s}}_{0}$ over states, the equation above fully determines the evolution of the probability over states with respect to time. If it is known that the state at time t1t-1 is $y$, then ${\bm{s}}_{t}$ is simply column $y$ of $\bm{P}$}. This implies that the sum of the elements of each column $y$ of $\bm{P}$} is one.
@@ D.2 Continuous-Time Markov Chains Over Finite-State Spaces
Given an initial probability distribution ${\bm{s}}_{0}$ over states, the equation above fully determines the evolution of the probability over states with respect to time. Indeed, the distribution over states at time $t$ is the solution of the linear ODE in (181): ${\bm{s}}_{t}=e^{t{\bm{Q}}}{\bm{s}}_{0}$. Therefore, if it is known that the state at time 0 is $j$, then ${\bm{s}}_{t}$ is simply column $j$ of $e^{t{\bm{Q}}}$. Naturally, one can rescale the time variable such that $t=\sigma(t)$, where $\sigma$ is monotonically increasing, $\sigma(0)=0$ and $\lim_{t\rightarrow 1}\sigma(t)=T$, so that for ${\bm{z}}_{0}:={\bm{s}}_{0}$, we get $e^{t{\bm{Q}}}{\bm{s}}_{0}=e^{\sigma(t){\bm{Q}}}{\bm{z}}_{0}={\bm{z}}_{t}$, and
@@
Matrices ${\bm{Q}}_{t}=\sigma^{{}^{\prime}}(t)({\bm{P}}-{\bm{I}})$ clearly satisfy the properties of transition-rate matrices (Suhov and Kelbert, 2008). Matrices ${\bm{Q}}_{t}$ are chosen such that: a) the matrix exponential $e^{\sigma(t){\bm{Q}}}$ is easy to calculate, which is essential as $p_{t|0}(x|y)=e^{\sigma(t){\bm{Q}}}(x,y)$; and b) ${\bm{z}}_{1}$ is an easy reference distribution to sample from (Austin et al., 2021; Campbell et al., 2022).